User name Peter L
Log entry time 09:12:50 on July23,2005
Entry number 149146
We started off doing a septum heating vs Right Septum Current. Placing the current at +10% lowered the right upper coil temp .08K limit and adding another 10% lowers it an additional .07K.
Septum current upper Coil T lower Coil T. Yoke T
275A 7.57 7.00 N/A
300A 7.50 6.97 N/A
Nominal 250A 7.65 7.07 5.41
275A Run 3774
300A Run 3775
Nominal Run 3776
Here is a list of run numbers and Septum temps corresponding to the rastor changes.
Run Right Upper Lower Yoke Rastor
3777 7.597 7.013 5.384 3X3
3778 7.647 7.050 5.420 5X5
3779 7.676 7.109 5.272 7X7
We saw that there was a change in the heating pattern due to rastor but it was not very drastic.
Next we changed the beam postition from the maximum of BPMB y=-2 to y=+3. Results are as follows:
Run BPMA BPMB RUpCoil RLoCoil RYoke Beam
3780 1.6,-1.2 -1.4,-1.0 6.97 6.365 5.02 9.94
3781 1.6,-1.4 -1.4,-2.0 6.97 6.365 5.02 9.94
3782 1.6,-.2 -1.4,1.0 6.95 6.351 5.016 9.94
3783 1.6,-.2 -1.4,2.0 6.96 6.355 5.018 9.94
3784 1.6,-.2 -1.4,3.0 7.02 6.39 5.04 9.94
Meanwhile Q2 tripped again and concluding this study I called a controlled access. I restarted Q2 and it ramped sucessfully until we called the access off and returned to the controil room. Now it states that there is an overtemp fault only instead of the Quench and Internal fault I got before.
Hachime also joined me in the hall to fix a compton problem, a restar of the ROC.
Throughout the evening we had a number of target scares cause by erratic CHL flows. The flow would simply decline from 23 to as low as 17 for no apparent reason. At this low amount of flow we lost control of our heating because the heater dropped to zero. Each time our remedy was to call MCC and have them turn off the beam. This is not at all a comfortable experience and we witnessed target pressures of higher than 245psi which is supposed to be venting.
TARGET OPERATORS PLEASE WATCH OUT FOR THIS. IF THE CHL FLOW DROPS AND YOU ARE RUNNING 35uA KILL THE BEAM BY CALLING MCC.